OpenStudy (anonymous):

a self employed contractor nearing retirement made two investments totaling $ 15 ,000.In one year .These investments yielded $ 1,023 in simple interest .part of the money was invested at 6% and the rest at 7.5 %.How much was invested at each rate ?

OpenStudy (anonymous):

let x and y represent your separate parts. x+y=15000 x=15000-y 1.06x+1.075y=1023+15000 1.06(15000-y)+1.075y=1023+15000 53/50(15000-y)+1.075y=1023+15000 300(53)-1.06y+1.075y=16023 .0175y=16023-15900 .0175Y=123 1/100+3/400=120 7/400)Y=120